# Boil-off of red supergiants: mass loss and type II-P supernovae

### v1

@@ 5.1 Comparison with red supergiant observations
Our model predicts that chromospheric material is clumpy, stochastically varying with time, and moving both inwards and outwards with typical velocities of $\sim$10 of km/s. These speeds are slightly smaller than observed for gas within 2 stellar radii of Antares and Betelgeuse, moving at velocities of $\sim\!20-50$ km/s (Ohnaka et al., 2017; López Ariste et al., 2022). We also expect that RSGs with larger atmospheric motions should have larger mass loss rates, as tentatively observed (Josselin & Plez, 2007). Such observations are typically made though near-infrared emission lines from molecules, which can extend out to a few stellar radii (Arroyo-Torres et al., 2015). It is not clear if our models can account for molecular emission at such large radii, and this should be investigated in future work.
@@
Our model predicts that most of RSG mass loss occurs through episodic ejection events. These mass ejection events would be accompanied by dust formation events, which may have caused the great dimming of Betelgeuse in 2020 (Montargès et al., 2021; Taniguchi et al., 2022). That event was preceded by radial velocity and molecular absorption disturbances in the photosphere several months before the dust formation (Dupree et al., 2022; Jadlovský et al., 2023), qualitatively consistent with our picture of the chromosphere responding to energetic convection below on a dynamical time scale of $\sim\!R_{\rm d}/v_{\rm esc}\sim 1\,{\rm yr}$. We discuss this further in Section 5.4.
@@ 5.2 Comparison with prior models
Although sometimes forgotten by the massive star community, models for shock-supported stellar chromospheres have been developed for decades (e.g., Hill & Willson 1979; Bertschinger & Chevalier 1985; Bowen 1988), primarily in the context of radially pulsating AGB stars and Mira variables. Our model is largely a reapplication of previous models to RSGs, but here the shocks are launched by convective motions rather than large-scale pulsation. This allows for analytic estimates of chromospheric densities and stellar mass loss rates, which can be readily compared with data or implemented into stellar evolution codes. The most important new aspect of our model is the inclusion of a spectrum of shock speeds $v_{s}$ rather than assuming a single value. The high-velocity tail of this spectrum greatly increases the time-averaged chromospheric density and mass loss rate.
Hill & Willson (1979) found approximately periodic and hydrostatic (in a time-averaged sense) chromospheric solutions as long as the dimensionless ratio $P_{0}/P$ (their equation 2) is small. In the relevant limit $v_{\rm con}\ll v_{\rm esc}$, this requires
@@ 5.3 Comparison with red supergiant simulations
It is useful to compare our model with simulations of RSGs, such as the 3D radiation hydrodynamical simulations of Goldberg et al. (2022a) (see also Chiavassa et al. 2011b; Chiavassa et al. 2011a; Ma et al. 2023; Chiavassa et al. 2024). These simulations show RMS near-surface convective velocities of $\sim$10 km/s, similar to those predicted by mixing length theory, for values of $\alpha_{\rm MLT}\sim 2-4$. These help justify our use MLT to compute the values of $v_{\rm con}$ in our models. However, our predicted mass loss rates are very sensitive to the convective velocities (see Section 5.6), so it will be necessary to directly compute mass loss in simulations in order to test and calibrate our model.
@@
Some very useful existing simulations are those of AGB stars and their dust-driven winds (Freytag & Höfner, 2008; Freytag et al., 2017; Freytag & Höfner, 2023). These simulations indeed exhibit RMS velocities of material in and above the photosphere that are nearly constant radius, with RMS radial velocities of $\sim$5-10 km/s and Mach numbers of order $\mathcal{M}\sim 2-3$. The density profiles of those models also resemble ours in Figure 2, with a sharp drop in density by a few orders of magnitude just above the surface of the star, followed by a more gradual decline at larger radii. Their AGB models have higher values of $v_{\rm con}/v_{\rm esc}$, cooler surface temperatures, and smaller dust formation radii than we expect for RSGs, so a detailed comparison should await simulations in the RSG regime.
@@ 5.5 Comparison with supernovae and SN 2023ixf
Another expectation of our model is that the CSM material is asymmetric and clumpy, stochastically varying in time. This is in line with observations of SN 2023ixf, where the early (and changing) continuum polarization indicates an aspherical CSM (Vasylyev et al., 2023). Spectral modeling of this event also suggests an aspherical CSM (Smith et al., 2023). We suspect an aspherical CSM will also lengthen the duration of shock breakout emission (Goldberg et al., 2022b) and rise to V-band maximum, which may help explain the slow rise of events like SN 2023ixf relative to models (e.g., Figure 9 and Hosseinzadeh et al. 2023). We note that type IIb SN 2013cu was inferred to have a smooth and spherical wind by modeling its narrow emission lines (Gräfener & Vink, 2016), but the hydrogen-poor progenitor of that event may have a different mass loss mechanism than our model for hydrogen-rich RSGs. Future hydrodynamic and radiative transfer modelling of our scenario will be needed for detailed predictions.
An interesting and unusual feature of SN 2023ixf is that its progenitor was clearly pulsating at large amplitude (Kilpatrick et al., 2023; Jencson et al., 2023), in contrast to most RSGs and SN progenitors. Such pulsation is thought to be crucial to the mass loss of AGB stars by driving shocks through the atmosphere (Höfner & Olofsson, 2018; Freytag et al., 2017). Our model could possibly account for this by replacing the convective velocity $v_{\rm con}$ with the pulsational velocity near the photosphere, which can be supersonic and hence larger than $v_{\rm con}$, driving higher mass loss rates. It is thus possible that the pulsations of SN 2023ixf’s progenitor created its unusually large CSM density and mass loss rate, as has been suggested previously (Yoon & Cantiello, 2010).
@@ 5.6 Caveats and uncertainties
One difference between these regimes is the convective Mach numbers, which are small in stellar interiors but are of order unity near the surface of a RSG. Convection is neither in the low-Mach number, incompressible regime of Kolmogorov turbulence, nor in the high-Mach number and highly compressible regime of Burgers turbulence. Three-dimensional simulations (e.g., Rabatin & Collins 2023a, b of such turbulence indicate that the convective speeds are fairly well approximated by a Maxwellian distribution, while the density is well approximated by a log-normal distribution, without strong correlations between density and velocity. However, this could be changed by effects of intermittency (Hopkins, 2013). Additionally, RSG atmospheres are not isotropic and can reach near-Eddington luminosities, greatly changing the nature of convection (e.g., Goldberg et al. 2022a) and correlations between density and velocity. Further numerical work on non-isotropic, moderate-Mach number, convectively driven turbulence is needed to address this uncertainty.
@@
We neglected radiative forces within the chromosphere, justified by the fact that typical opacities at these temperatures are of order $\kappa\sim 10^{-4}-10^{-2}\,{\rm cm^{2}}/{\rm g}$ (Plez et al., 1992). The corresponding Eddington factor, i.e., the ratio between radiative acceleration and gravity, is $f_{\rm ed}=g/g_{\rm rad}=L/L_{\rm ed}=L\kappa/(4\pi GMc)\sim 10^{-4}-10^{-2}$ at these opacities. This is approximately consistent with radiative transfer modeling of red giant atmospheres (Ireland et al., 2008) that find only small radiative accelerations within the chromosphere, although it has been suggested turbulent motions may decrease line shadowing and increase the effective opacity enough to have a substantial effect (Josselin & Plez, 2007). Since our predicted mass loss rates are strongly dependent on $v_{\rm esc}$ and hence on the net effect of gravity and radiation forces, even small radiative forces could substantially increase chromospheric densities and associated mass loss rates, and should be incorporated in more sophisticated versions of our model.
@@
We also neglected the effects of rotation and magnetic fields in our models, which is reasonable given the small rotation rates of RSGs, and the lack of strong surface fields. The cool chromospheric material is almost completely neutral, diminishing the importance of magnetic fields. Furthermore, the thermal and turbulent pressures of our time-averaged model shown in Figure 2 are larger than the magnetic pressure of a dipole field for a surface field strength of 1 G, so fairly large magnetic fields are required in order to dominate the dynamics. Thirumalai & Heyl (2010, 2012) investigated a hybrid magnetic Weber-Davis and dust-driven wind model for red giants and RSGs and its affect on the terminal wind speeds. However, the mass loss rate is imposed in such models (rather than being calculated from first principles), so such models do not solve the wind launching problem below the dust formation radius.
@@
Finally, our model for the dust-driven wind above the dust formation radius is undoubtedly too simple. More realistic dust chemistry, grain formation, and radiative driving should be developed for a better model of the wind material. Additionally, resonant drag instabilities will occur in the dust-driven ouflow (Squire & Hopkins, 2018; Moseley et al., 2019; Steinwandel et al., 2022), which may affect the outflow speed and appearance. Understanding such phenomena may be needed not only for mass loss models, but also for more accurate observational mass loss rate estimates.
@@ 6 Conclusions
We have developed a simple model for material supported by shock waves in the chromospheres of red supergiants (RSGs). These chromospheres extend roughly from the stellar surface to the dust formation radius at $\sim$5-10 stellar radii. Gravity is balanced by momentum deposition from outgoing shock waves, whose effect is similar to turbulent pressure support. For our assumption of nearly constant shock speeds (justified by simulation results), the density profile is similar to that of an isothermal atmosphere with scale height $v_{\rm con}^{2}/g\ll R$, which would cause a steeply decreasing density for shock speeds comparable to the surface convective speeds $v_{\rm con}$. However, we also showed that averaging over a stochastically varying convective velocity spectrum greatly increases the chromospheric density, causing it to fall off less steeply with radius.
@@
7 Acknowledments
We thank Luc Dessart and Jared Goldberg for useful discussions, and Kostas Antoniadis for supplying mass loss measurements. This work benefited from collaborative meetings supported by the Gordon and Betty Moore Foundation (grant GBMF5076). This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. This research is also supported by the Sherman Fairchild Postdoctoral Fellowship at the California Institute of Technology.

### v2

@@ 5.1 Comparison with red supergiant observations
Our model predicts that chromospheric material is clumpy, stochastically varying with time, and moving both inwards and outwards with typical velocities of $\sim$10 of km/s. These speeds are slightly smaller than observed for gas within 2 stellar radii of Antares and Betelgeuse, moving at velocities of $\sim\!20-50$ km/s (Ohnaka et al., 2017; López Ariste et al., 2022). We also expect that RSGs with larger atmospheric motions should have larger mass loss rates, as tentatively observed (Josselin and Plez, 2007). Such observations are typically made though near-infrared emission lines from molecules, which can extend out to a few stellar radii (Arroyo-Torres et al., 2015). It is not clear if our models can account for molecular emission at such large radii, and this should be investigated in future work.
@@
Our model predicts that most of RSG mass loss occurs through episodic ejection events driven by shocks. These mass ejection events would be accompanied by dust formation events, which may have caused the great dimming of Betelgeuse in 2020 (Montargès et al., 2021; Taniguchi et al., 2022). That event was preceded by radial velocity and molecular absorption disturbances in the photosphere several months before the dust formation (Dupree et al., 2022; Jadlovský et al., 2023). Additional tomography of Betelgeuse showed two strong outgoing shock waves around the time of the great dimming event (Kravchenko et al., 2021; Jadlovský et al., 2024). These observations are qualitatively consistent with our picture of the chromosphere expanding and forming dust during mass ejection events, on a time scale of $\sim\!R_{\rm d}/v_{\rm esc}\sim 1\,{\rm yr}$. We discuss this further in Section 5.4.
@@ 5.2 Comparison with prior models
Although sometimes forgotten by the massive star community, models for shock-supported stellar chromospheres have been developed for decades (e.g., Hill and Willson 1979; Bertschinger and Chevalier 1985; Bowen 1988), primarily in the context of radially pulsating AGB stars and Mira variables. Our model is largely a reapplication of previous models to RSGs, but here the shocks are launched by convective motions rather than large-scale pulsation. This allows for analytic estimates of chromospheric densities and stellar mass loss rates, which can be readily compared with data or implemented into stellar evolution codes. The most important new aspect of our model is the inclusion of a spectrum of shock speeds $v_{s}$ rather than assuming a single value. The high-velocity tail of this spectrum greatly increases the time-averaged chromospheric density and mass loss rate.
Hill and Willson (1979) found approximately periodic and hydrostatic (in a time-averaged sense) chromospheric solutions as long as the dimensionless ratio $P_{0}/P$ (their equation 2) is small. In the relevant limit $v_{\rm con}\ll v_{\rm esc}$, this requires
@@ 5.3 Comparison with red supergiant simulations
It is useful to compare our model with simulations of RSGs, such as the 3D radiation hydrodynamical simulations of Goldberg et al. (2022a) (see also Chiavassa et al. 2011a, b; Ma et al. 2023; Chiavassa et al. 2024). These simulations show RMS near-surface convective velocities of $\sim$10 km/s, similar to those predicted by mixing length theory, for values of $\alpha_{\rm MLT}\sim 2-4$. These help justify our use MLT to compute the values of $v_{\rm con}$ in our models. However, our predicted mass loss rates are very sensitive to the convective velocities (see Section 5.6), so it will be necessary to directly compute mass loss in simulations in order to test and calibrate our model.
@@
Some very useful existing simulations are those of AGB stars and their dust-driven winds (Freytag and Höfner, 2008; Freytag et al., 2017; Freytag and Höfner, 2023). These simulations indeed exhibit RMS velocities of material in and above the photosphere that are nearly constant radius, with RMS radial velocities of $\sim$5-10 km/s and Mach numbers of order $\mathcal{M}\sim 2-3$. The density profiles of those models also resemble ours in Figure 2, with a sharp drop in density by a few orders of magnitude just above the surface of the star, followed by a more gradual decline at larger radii. Their AGB models have higher values of $v_{\rm con}/v_{\rm esc}$, cooler surface temperatures, and smaller dust formation radii than we expect for RSGs, so a detailed comparison should await simulations in the RSG regime.
@@ 5.5 Comparison with supernovae and SN 2023ixf
Another expectation of our model is that the CSM material is asymmetric and clumpy, stochastically varying in time. This is in line with observations of SN 2023ixf, where the early (and changing) continuum polarization indicates an aspherical CSM (Vasylyev et al., 2023). Spectral modeling of this event also suggests an aspherical CSM (Smith et al., 2023). We suspect an aspherical CSM will also lengthen the duration of shock breakout emission (Goldberg et al., 2022b) and rise to V-band maximum, which may help explain the slow rise of events like SN 2023ixf relative to models (e.g., Figure 9 and Hosseinzadeh et al. 2023). We note that type IIb SN 2013cu was inferred to have a smooth and spherical wind by modeling its narrow emission lines (Gräfener and Vink, 2016), but the hydrogen-poor progenitor of that event may have a different mass loss mechanism than our model for hydrogen-rich RSGs. Future hydrodynamic and radiative transfer modelling of our scenario will be needed for detailed predictions.
An interesting and unusual feature of SN 2023ixf is that its progenitor was clearly pulsating at large amplitude (Kilpatrick et al., 2023; Jencson et al., 2023). Many RSGs in the Milky Way and M31 are also observed to pulsate fairly coherently, though often at somewhat smaller amplitudes (Soraisam et al., 2018). Similar pulsations are thought to be crucial to the mass loss of AGB stars by driving shocks through the atmosphere (Höfner and Olofsson, 2018; Freytag et al., 2017). Our model could possibly account for this by replacing the convective velocity $v_{\rm con}$ with the pulsational velocity near the photosphere, which can be supersonic and hence larger than $v_{\rm con}$, driving higher mass loss rates. It is thus possible that the pulsations of SN 2023ixf’s progenitor created its unusually large CSM density and mass loss rate, as has been suggested previously (Yoon and Cantiello, 2010).
@@ 5.6 Caveats and uncertainties
One difference between these regimes is the convective Mach numbers, which are small in stellar interiors but are of order unity near the surface of a RSG. Convection is neither in the low-Mach number, incompressible regime of Kolmogorov turbulence, nor in the high-Mach number and highly compressible regime of Burgers turbulence. Three-dimensional simulations (e.g., Rabatin and Collins 2023a, b of such turbulence indicate that the convective speeds are fairly well approximated by a Maxwellian distribution, while the density is well approximated by a log-normal distribution, without strong correlations between density and velocity. However, this could be changed by effects of intermittency (Hopkins, 2013). Additionally, RSG atmospheres are not isotropic and can reach near-Eddington luminosities, greatly changing the nature of convection (e.g., Goldberg et al. 2022a) and correlations between density and velocity. Further numerical work on non-isotropic, moderate-Mach number, convectively driven turbulence is needed to address this uncertainty.
@@
We neglected radiative forces within the chromosphere, justified by the fact that typical opacities at these temperatures are of order $\kappa\sim 10^{-4}-10^{-2}\,{\rm cm^{2}}/{\rm g}$ (Plez et al., 1992). The corresponding Eddington factor, i.e., the ratio between radiative acceleration and gravity, is $f_{\rm ed}=g/g_{\rm rad}=L/L_{\rm ed}=L\kappa/(4\pi GMc)\sim 10^{-4}-10^{-2}$ at these opacities. This is approximately consistent with radiative transfer modeling of red giant atmospheres (Ireland et al., 2008) that find only small radiative accelerations within the chromosphere, although it has been suggested turbulent motions may decrease line shadowing and increase the effective opacity enough to have a substantial effect (Josselin and Plez, 2007). Since our predicted mass loss rates are strongly dependent on $v_{\rm esc}$ and hence on the net effect of gravity and radiation forces, even small radiative forces could substantially increase chromospheric densities and associated mass loss rates, and should be incorporated in more sophisticated versions of our model.
@@
We also neglected the effects of rotation and magnetic fields in our models, which is reasonable given the small rotation rates of RSGs, and the lack of strong surface fields. The cool chromospheric material is almost completely neutral, diminishing the importance of magnetic fields. Furthermore, the thermal and turbulent pressures of our time-averaged model shown in Figure 2 are larger than the magnetic pressure of a dipole field for a surface field strength of 1 G, so fairly large magnetic fields are required in order to dominate the dynamics. Thirumalai and Heyl (2010, 2012) investigated a hybrid magnetic Weber-Davis and dust-driven wind model for red giants and RSGs and its affect on the terminal wind speeds. However, the mass loss rate is imposed in such models (rather than being calculated from first principles), so such models do not solve the wind launching problem below the dust formation radius.
@@
Finally, our model for the dust-driven wind above the dust formation radius is undoubtedly too simple. More realistic dust chemistry, grain formation, and radiative driving should be developed for a better model of the wind material. Additionally, resonant drag instabilities will occur in the dust-driven ouflow (Squire and Hopkins, 2018; Moseley et al., 2019; Steinwandel et al., 2022), which may affect the outflow speed and appearance. Understanding such phenomena may be needed not only for mass loss models, but also for more accurate observational mass loss rate estimates.
@@ 6 Conclusions
We have developed a simple model for material supported by shock waves in the chromospheres of red supergiants (RSGs). These chromospheres extend roughly from the stellar surface to the dust formation radius at $\sim$5 stellar radii. Gravity is balanced by momentum deposition from outgoing shock waves, whose effect is similar to turbulent pressure support. For our assumption of nearly constant shock speeds (justified by simulation results), the density profile is similar to that of an isothermal atmosphere with scale height $v_{\rm con}^{2}/g\ll R$, which would cause a steeply decreasing density for shock speeds comparable to the surface convective speeds $v_{\rm con}$. However, we also showed that averaging over a stochastically varying convective velocity spectrum greatly increases the chromospheric density, causing it to fall off less steeply with radius.
@@
Acknowledgments
We thank Luc Dessart, Jared Goldberg, Shazrene Mohamed, Konstantin Batygin, and Lars Bildsten for useful discussions, and Kostas Antoniadis for supplying mass loss measurements. This work benefited from collaborative meetings supported by the Gordon and Betty Moore Foundation (grant GBMF5076). This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. This research is also supported by the Sherman Fairchild Postdoctoral Fellowship at the California Institute of Technology.